# Science Prospects for the Southern Wide-field Gamma-ray Observatory: SWGO

### v1

@@ 3.5 Outer Array Options
Some full-scale prototypes have been successfully produced and shipped to Chile for SWGO-PF (see Figure 5), including one with a reflective nano-coating and one that is in operation equipped with a multi-PMT module. This module is composed of 73” PMTs to maximize the angular acceptance of Cherenkov light produced by air-shower particles that penetrate the tank, and to provide directionality information at each WCD unit.
@@ 3.6 The Radio Array SWGO-TURBO
An array of 800 radio antennas, named The Utility for Radio Beam-formed Observations (TURBO), will be constructed as part of the SWGO Project, following a successful bid to the ERC Consolidator Grants. The dual-polarized antennas will observe the pulsed radio emission from air showers in the 50-200 MHz frequency band, and data acquisition will be fully integrated within the observatory. Since SWGO-TURBO is already funded, it will be deployed during the early phase(s) of SWGO and therefore complement the particle observations for the full operational time of the observatory. The readout of the radio array will be initiated by a trigger from the particle detectors. By applying interferometric reconstruction, it is anticipated that the influence of noise can be significantly reduced compared to traditional radio observations of air showers.
@@ 5.2.1 Galactic Diffuse Emission
While HAWC (Alfaro et al., 2024a) and LHAASO (Cao et al., 2025) have deeply surveyed the northern sky, including measurements of the diffuse emission, SWGO will extend the survey to the whole southern sky with improved angular and energy resolution. SWGO’s superior gamma-hadron separation provides better sensitivity than current instruments.
@@ 5.3 Ultra-High-Energy Sources and PeVatrons
The search for PeVatrons is directly connected to the long-standing mystery of the origin of CRs. The CR spectrum follows a PL distribution with an index of -2.7 above 30 GeV, transitioning at 3 PeV to a steeper slope of -3.0, forming the well-known “knee” feature (e.g. Blümer et al., 2009). Despite advances in measurements of CRs arriving at Earth up to and beyond the knee, it remains unclear which acceleration sites produce them. This limitation is well known, as particles are randomly scattering in the galactic magnetic field before arriving at Earth. In the PeV regime, it is thus necessary to look for neutral messengers such as UHE gamma-ray and neutrinos. Despite recent groundbreaking advancements in identifying PeVatrons, no galactic source has yet been unambiguously confirmed as an accelerator of hadrons to PeV energies, largely due to the difficulty in disentangling hadronic and leptonic emission scenarios. Recent studies suggest that UHE gamma-ray emission may be a universal feature of the environments of powerful pulsars, particularly those with high spin-down luminosity (e.g. Albert et al., 2021; Breuhaus et al., 2021; de Oña Wilhelmi et al., 2022), indicating that the UHE source population is dominated by PWNe. Since inverse-Compton scattering of PeV electrons on the CMB can readily produce UHE gamma rays, detecting UHE emission up to 1 PeV alone cannot conclusively establish a hadronic origin (see for example [)2021ApJ…908L..49B. Consequently, multiwavelength (MWL) and multi-messenger approaches become crucial for resolving this ambiguity. A clear correlation with dense target material, such as a MC, would support a hadronic scenario. However, an undeniable confirmation of hadronic PeVatron activity would come from the simultaneous co-detection of neutrinos and UHE gamma rays.
@@ 8.1 Spectrum and Composition of CRs around the Knee
The mechanism behind the acceleration of CRs and the physical processes that affect their propagation through space are subjects of intense debate. An indirect approach to investigating these is to study the CR energy spectrum and mass composition. In the energy region around and below the knee (10 TeV 10 PeV), where SWGO will be sensitive, the energy spectra of the elemental mass groups of CRs have not been thoroughly studied, and several related questions need to be solved (Mollerach & Roulet, 2018; Piazzoli et al., 2022). The first question is related to the existence of a softening feature in the spectrum of Fe nuclei at hundreds of TeV. Results of the NUCLEON satellite implied the presence of a Peter’s cycle (Peters, 1961) in the 10 TeV - 1 PeV interval and, in particular, individual cutoffs in the elemental spectra of light and heavy nuclei at rigidities (Atkin et al., 2017). DAMPE (An et al., 2019; Alemanno et al., 2021), CALET (Adriani et al., 2019, 2023) and updated data from NUCLEON (Grebenyuk et al., 2019) have confirmed the existence of cutoffs in the CR spectra of light elements: close to 14 TeV for protons (or H nuclei) and around 34 TeV for He nuclei, which is in agreement with the above picture. Accordingly, cutoffs in the CR spectra of the C-N-O and Fe mass groups should be expected between 90 TeV and 1 PeV. In particular, the breaks for the C-N-O and Fe spectra must be located at around 90 TeV and 300 TeV, respectively. The measurement of these features will be crucial to confirm the Peter’s cycle scenario below the knee and to test different hypotheses about the origin of CRs in this energy regime (Liu et al., 2019; Yue et al., 2020; Malkov & Moskalenko, 2022).
Another open question appears around the knee itself. Here, measurements of the air-shower experiments KASCADE (Antoni et al., 2005; Apel et al., 2009, 2013), IceTop / IceCube (Aartsen et al., 2019), MAKET-ANI (Chilingarian et al., 2004, 2007), and EAS-TOP / MACRO (Aglietta et al., 2004) point out that this softening is produced by cutoffs in the spectra of light primaries (H and He nuclei) due to another Peter’s cycle in the 1-100 PeV energy region (Apel et al., 2013). However, results from the ARGO-YBJ (Bartoli et al., 2015; Montini et al., 2016) and Tibet-ASgamma (Amenomori et al., 2019) experiments indicate that the cutoff in the spectrum of the H+He mass group lies at around 400-700 TeV instead of PeV energies. These findings are however in conflict with the latest LHAASO results (The LHAASO Collaboration et al., 2025), showing that there is much that is still to be understood. It remains an open question if the knee is the point where CRs start to escape from their magnetic confinement at the sources (Cristofari et al., 2020; Vink, 2022) or in our galaxy (Hörandel, 2004; Candia et al., 2002; Giacinti & Semikoz, 2023). To better answer these questions, further measurements of the spectra and mass composition of CRs around and below the knee with larger statistics and improved precision are required.
@@
In addition, enhanced muon-counting capabilities will improve the determination of primary CR composition. Generally, identifying the species of a CR primary using WCDs is challenging, as the longitudinal profile of the shower is not accessible. However, the dual-layer tank design makes the larger top layer act as a calorimeter for the electromagnetic component of the shower while allowing the muons to reach the bottom layer, making it a crucial tool for identifying and, ideally, counting them. The relation between the total electromagnetic energy, , and the total number of muons, , of an air shower may be used as a strong separator between different CR species. See Section 8.3 for more details regarding SWGO’s study of the muon lateral distributions.
@@ 8.2 Anisotropy
The study of large-scale anisotropies can lead to strong biases if only partial FoVs are considered. Ground-based instruments have access only to a fraction of the sky, and complementary measurements from different experiments are desirable for achieving full-sky coverage. This was first done in the TeV region by HAWC and IceCube (Abeysekara et al., 2019). However, combining data remains challenging due to differences in detection techniques as well as sensitivities to energy and composition across experiments. Any attempt to align the systematic uncertainties is further complicated by the small region of overlap ( for the HAWC and IceCube measurements) in which the direction reconstruction quality is the poorest for both experiments. SWGO will contribute significantly to studying the full-sky dipole anisotropy by providing large statistics and high-quality data, employing a detection mechanism similar to that of HAWC and a larger overlap with its FoV. As shown in Figure 27, even a simple plane fit reconstruction for the arrival directions can help achieve angular resolutions , enough to explore mid-scale anisotropies, for which the size of structures is of the order of tens of degrees.

### v2

@@ 3.5 Outer Array Options
Some full-scale prototypes have been successfully produced and shipped to Chile for SWGO-PF (see Figure 5), including one with a reflective nano-coating and one that is in operation equipped with a multi-PMT module (Figure 6). This module is composed of 73” PMTs to maximize the angular acceptance of Cherenkov light produced by air-shower particles that penetrate the tank, and to provide directionality information at each WCD unit.
@@ 3.6 The Radio Array SWGO-TURBO
An array of 800 radio antennas, named The Utility for Radio Beam-formed Observations (TURBO), will be constructed as part of the SWGO Project, following a successful bid to the ERC Consolidator Grants. The dual-polarized antennas will observe the pulsed radio emission from air showers in the 50-200 MHz frequency band, and data acquisition will be fully integrated within the observatory. Since SWGO-TURBO is already funded, it will be deployed during the early phase(s) of SWGO and therefore complement the particle observations for the full operational time of the observatory. The readout of the radio array will be initiated by a trigger from the particle detectors. By applying interferometric reconstruction Schoorlemmer & Carvalho (2021); Schoorlemmer, H [Pierre Auger Collaboration] (2023), it is anticipated that the influence of noise can be significantly reduced compared to traditional radio observations of air showers.
@@ 5.2.1 Galactic Diffuse Emission
While HAWC (Alfaro et al., 2024a) and LHAASO (Cao et al., 2025a) have deeply surveyed the northern sky, including measurements of the diffuse emission, SWGO will extend the survey to the whole southern sky with improved angular and energy resolution. SWGO’s superior gamma-hadron separation provides better sensitivity than current instruments.
@@ 5.3 Ultra-High-Energy Sources and PeVatrons
The search for PeVatrons is directly connected to the long-standing mystery of the origin of CRs. The CR spectrum follows a PL distribution with an index of -2.7 above 30 GeV, transitioning at 3 PeV to a steeper slope of -3.0, forming the well-known “knee” feature (e.g. Blümer et al., 2009). Despite advances in measurements of CRs arriving at Earth up to and beyond the knee, it remains unclear which acceleration sites produce them. This limitation is well known, as particles are randomly scattering in the galactic magnetic field before arriving at Earth. In the PeV regime, it is thus necessary to look for neutral messengers such as UHE gamma-ray and neutrinos. Despite recent groundbreaking advancements in identifying PeVatrons, no galactic source has yet been unambiguously confirmed as an accelerator of hadrons to PeV energies, largely due to the difficulty in disentangling hadronic and leptonic emission scenarios. Recent studies suggest that UHE gamma-ray emission may be a universal feature of the environments of powerful pulsars, particularly those with high spin-down luminosity (e.g. Albert et al., 2021; Breuhaus et al., 2021; de Oña Wilhelmi et al., 2022), indicating that the UHE source population is dominated by PWNe. Since inverse-Compton scattering of PeV electrons on the CMB can readily produce UHE gamma rays, detecting UHE emission up to 1 PeV alone cannot conclusively establish a hadronic origin (see for example Breuhaus et al., 2021). Consequently, multiwavelength (MWL) and multi-messenger approaches become crucial for resolving this ambiguity. A clear correlation with dense target material, such as a MC, would support a hadronic scenario. However, an undeniable confirmation of hadronic PeVatron activity would come from the simultaneous co-detection of neutrinos and UHE gamma rays.
@@ 8 Cosmic-Ray (CR) Measurements
In addition to a vast range of gamma-ray topics, SWGO will act as a unique tool for studying charged cosmic rays in the TeV-PeV regime. Its large effective area and advanced muon counting capability, combined with its privileged location in the Southern Hemisphere, will allow the pursuit of relevant open questions in TeV-PeV cosmic-ray astrophysics. The main topics of interest are described in detail below.
@@ 8.1 Spectrum and Composition of CRs around the Knee
The mechanism behind the acceleration of CRs and the physical processes that affect their propagation through space are subjects of intense debate. An indirect approach to investigating these is to study the CR energy spectrum and mass composition. In the energy region around and below the knee (1 TeV 10 PeV), where SWGO will be sensitive, there have recently been several experimental efforts to study the energy spectrum the bulk of cosmic rays and the relative abundances elemental mass groups of this radiation. These aim to fill the gap between direct and indirect CR data and to refine the measurements around the knee to investigate the properties of galactic CRs. Between TeV and PeV, measurements on the all-particle spectrum of CRs from the NUCLEON satellite (Atkin et al., 2017; Grebenyuk et al., 2019) and the HAWC observatory (Alfaro et al., 2017, 2025) have established a first bridge with high statistics and large precision between the TeV CR balloon- and space-borne experiments and the PeV air-shower observations and have showed the existence of a softening at tens of TeV. The last analysis of HAWC has located this structure at TeV (Alfaro et al., 2025). More recently, LHAASO has measured the total spectrum around the knee, from TeV to PeV, with a precision which have not been achieved so far by other experiments (The LHAASO Collaboration et al., 2025). The position of the knee was found at PeV. LHAASO spectrum is in agreement with HAWC data between TeV and 1 PeV and shows no features below the knee.
Measurements of the elemental mass groups of CRs in the TeV energy region have also recently discovered several breaks in their energy spectra. NUCLEON (Grebenyuk et al., 2019), DAMPE (An et al., 2019; Alemanno et al., 2021) and CALET (Adriani et al., 2019, 2023) have observed the existence of cutoffs in the CR spectra of light elements: close to TeV for protons (or H nuclei) and around TeV for He primaries. DAMPE Alemanno et al. (2024) and HAWC Albert et al. (2022) have also observed a softening in the spectrum of HHe nuclei at . Preliminary results from HAWC point towards the TeV softening of the all-particle CR spectrum is due to the cutoffs in the H and He components of CRs and an increment in the contribution of heavy primaries Arteaga-Velázquez et al. (2024). Hardenings at around in the spectra of H and He nuclei have been also reported by HAWC Arteaga-Velázquez et al. (2024). GRAPES-3 has confirmed such feature for the spectrum of protons at TeV Varsi et al. (2024). LHAASO has provided recent data on the proton spectrum with the largest precision to date within the energy range and has detected a hardening close to TeV (Cao et al., 2025b). It should be pointed out that proton flux measured by LHAASO is larger than that measured by GRAPES-III and HAWC between TeV and PeV. Regarding heavier CR primaries, NUCLEON results have hinted at the presence of individual cutoffs in the elemental spectra of heavy nuclei at rigidities (Atkin et al., 2017). HAWC has also pointed out a softening in the spectrum of the heavy (C-Fe) component of CRs at hundreds of TeV Arteaga-Velázquez et al. (2024).
At energies around the knee, measurements of the air-shower experiments KASCADE (Antoni et al., 2005; Apel et al., 2009, 2013), IceTop / IceCube (Aartsen et al., 2019) and MAKET-ANI (Chilingarian et al., 2004, 2007) indicate that this softening is produced by cutoffs in the spectra of light primaries (H and He nuclei). Also, the latest results of LHAASO on the mean logarithmic mass supports the picture that the knee is dominated by light primaries (The LHAASO Collaboration et al., 2025), and the measurements on the energy spectrum of protons confirm the existence of a cutoff at energies around the knne ( PeV) (Cao et al., 2025b). However, results from the ARGO-YBJ (Bartoli et al., 2015; Montini et al., 2016) and Tibet-ASgamma (Amenomori et al., 2019) experiments show a cutoff in the spectrum of the H+He mass group lies at around 400-700 TeV. More accurate measurements of the elemental spectra over the full energy range are needed to determine whether this is an additional feature or whether this is in tension with the other experimental results of the knee being due to a softening of light CRs. KASCADE-Grande data have also shown individual cutoffs in the spectra of the CNO and Fe components. Interestingly, within systematic and statistical errors, the positions of the breaks detected by KASCADE-Grande in the elemental groups of CRs seem to depend on the charge of the primary particle (Antoni et al., 2005; Apel et al., 2009, 2013).
The results of KASCADE-Grande on the knee-like features in the spectra of the mass groups point to a rigidity dependence consistent with a Peters cycle (Peters, 1961), as the maximum acceleration energy of a source population is expected to be proportional to the nuclear charge. However, also propagation effects on the cosmic-ray spectra are expected to depend on the rigidity, and the knee would then indicate the point where galactic PeV CRs start to escape from their magnetic confinement at the sources (Hörandel, 2004; Cristofari et al., 2020; Vink, 2022) and/or in the Milky Way (Candia et al., 2002; Hörandel, 2004; Giacinti & Semikoz, 2023). The origin of the TeV knee-like feature in the all-particle spectrum of CRs may have a similar origin according to NUCLEON results. If confirmed, this scenario would predict charge-dependent cutoffs in the TeV CR spectra of the heavier nuclei, too. For the CNO and Fe mass groups, the cutoffs would lie between and TeV. The existence of these features could indicate the presence of an individual nearby TeV CR source inside our galaxy (Liu et al., 2019) or a complete population of such a kind of CR sources (Yue et al., 2020). It could also point out an unknown mechanism affecting the propagation of TeV CR in the Milky Way (Malkov & Moskalenko, 2022). The hardenings detected at around TeV by HAWC Arteaga-Velázquez et al. (2024) and Grapes-3 Varsi et al. (2024) in the spectra of the light CR nuclei may indicate the point where the contribution from these phenomena is surpassed by the contribution from the background of PeV CR sources.
In order to test different hypotheses about the origin of CRs at TeV and PeV energies, the shape of the energy spectra and the location of the knee-like features for the elemental mass groups of CRs must be measured with high precision. SWGO can contribute to this task by measuring the elemental spectra in the TeV energy regime for H, He, CNO and Fe primary nuclei. The H and He TeV cutoffs need confirmation with large statistics and high precision, as well as the CNO and Fe TeV softenings, which have just been hinted at by some experiments. SWGO data would also be important in investigating the transition from the TeV to the PeV energy regime in the spectra of CRs by measuring the TeV hardenings in the spectra of CRs with high precision. In addition, SWGO can contribute to measuring the precise location of the p, He, and CNO knee-like features at PeV energies to test models about the production and propagation of galactic CRs. LHAASO results on the energy spectrum of protons and on the abundance of light primaries around the knee need to be confirmed, as well as the HAWC results on the origin of the TeV cutoff at the total CR spectrum. It is important to point out that SWGO measurements in the TeV energy range will also be important to cross-check the energy scales of current direct and indirect experiments.
@@
In addition, enhanced muon-counting capabilities will improve the determination of the primary CR composition. Generally, identifying the species of a CR primary using WCDs is challenging, as the longitudinal profile of the shower is not measured directly. However, the dual-layer tank design makes the larger top layer act as a calorimeter for the electromagnetic component of the shower while allowing the muons to reach the bottom layer, making it a crucial tool for identifying and, ideally, counting them. The relation between the total electromagnetic energy, , and the total number of muons, , of an air shower may be used as a strong separator between different CR species. See Section 8.3 for more details regarding SWGO’s study of the muon lateral distributions.
@@ 8.2 Anisotropy
The study of large-scale anisotropies can lead to strong biases if only partial FoVs are considered. Ground-based instruments have access only to a fraction of the sky, and complementary measurements from different experiments are desirable for achieving full-sky coverage. This was first done in the TeV region by HAWC and IceCube (Abeysekara et al., 2019). However, combining data remains challenging due to differences in detection techniques as well as sensitivities to energy and composition across experiments. Any attempt to align the systematic uncertainties is further complicated by the small region of overlap ( for the HAWC and IceCube measurements) in which the direction reconstruction quality is the poorest for both experiments. SWGO will contribute significantly to studying the full-sky dipole anisotropy by providing large statistics and high-quality data, employing a detection mechanism similar to that of HAWC and a larger overlap with its FoV. For all considered primaries, even a simple plane fit reconstruction for the arrival directions can help achieve angular resolutions , enough to explore mid-scale anisotropies, for which the size of structures is of the order of tens of degrees.